OpenStudy (anonymous):

A Bag contains 3 green marble and 8 whites marble. suppose a marble is randomly selected. what are the odds in favor of picking a white marble. A.8 to 3 B. 8 to 11 C. 3 to 8 D. 3 to 11

Directrix (directrix):

A.8 to 3 (Chances for) P(x) = --------------- (Total chances) So, for example, the probability of drawing an ace in a single deck of 52 cards is 4/52 = 1/13 (or about 0.077 = 7.7%). Odds, on the other hand, are given as: (Chances for) : (Chances against) http://mathforum.org/library/drmath/view/56706.html